OpenStudy (dcolley):

The minute hand of a clock is 10 inches long and moves from 12 to 11o'clock. How far does the tip of the minute hand move? Express your answer in terms of π and then round two decimal places.

OpenStudy (anonymous):

circumference of that sircle is 2*pi*10 = 20pi from 12 to 11 o'clock is 11/12ths of that circumference 11/12 * 20 = 18 1/3 pi = 57.60

OpenStudy (dcolley):

Thanks!! I kept getting the first part of the question wrong

OpenStudy (phi):